OpenStudy (anonymous):

What percent of 60 is 12? Show your work.

OpenStudy (anonymous):

\[\LARGE 60\cdot x\% =12\] \[\LARGE 60\cdot \frac{x}{100}=12\] \[\LARGE \frac{60x}{100}=12\] \[\LARGE \frac{6x}{10}=12\] \[\LARGE 6x=10\cdot 12\] can you do it now ? :)

OpenStudy (anonymous):

the unknown in this problem is the rate or percentage

OpenStudy (anonymous):

i think he can do it now

OpenStudy (anonymous):

when i get 1.2 i times it by 6 right

OpenStudy (anonymous):

you shouldn't get 1.2 ! it's 12

OpenStudy (anonymous):

auff... it's 20 LOL ahhaha...

OpenStudy (anonymous):

oooh ok i did the % of 12 and thats 1.2

OpenStudy (anonymous):

you can do it :)

OpenStudy (anonymous):

i ment 120

OpenStudy (anonymous):

\[\LARGE 6x=120 \quad \Longrightarrow \quad x=\frac{120}{6}\] x=20

OpenStudy (anonymous):

@didierdre I don't believe this is the number we're looking for !! can you prove to me that it is ??

OpenStudy (anonymous):

so do you always * by 10

OpenStudy (anonymous):

@Kreshnik how i suppose to do that

OpenStudy (anonymous):

6*20=120

OpenStudy (anonymous):

If you read the question again it says: What percent of 60 is 12? now we were looking for this percent... Let's guess we've found it . And it's 20% substitute 20 in x... I'll write down equation for you. \[\LARGE 60\cdot \frac{x}{100}=12\] substitute x=20 you should get both sides equal . :)

OpenStudy (anonymous):

12/60 x 100 = 20 percent

OpenStudy (anonymous):

does x means "times" ... \[\LARGE \frac{12}{60}\cdot 100= 12 \quad \quad ?\]